the equilibrium constant for the reaction below decreases from 20570.0 at 440.0 C to 23.0 at 900.0 c
4NO + 3O2 .....2N2O5
from this data...calculate the value of change of heat for the reaction
Answer #1


The relation between the equilibrium constant and heat for the reaction is given by Van't Hoff equation

ln(K2/K1) = H/R[1/T1-1/T2] or

log(K2/K1) = (H/(2.303*R)*(1/T1-1/T2)

K2 = 23    T2 = 1173 K

K1 = 20570    T1 = 713 K

H = value of change of heat for the reaction

log(23/20570) = (H/(2.303*8.314)*(1/713-1/1173)

The value of change of heat for the reaction, Hrxn = -102.748 KJ